# Complete Synchronization and Pattern Selection through Amplitude Dynamics and Diffusion in Heterogeneous Oscillatory Media

### v1

@@ 3.2 Comparison of experimental and numerical results
Despite this agreement, systematic differences remain. Experimentally, high-amplitude regions pin to edges and form anisotropic shapes, while in simulations, they nucleate as compact patches. Such discrepancies arise from inhomogeneities, gradients, and boundary effects absent in the model, which assumes randomly distributed parameters (frequencies) and Neumann boundary conditions. Quantitative amplitude, phase, and frequency deviations are also natural, since the SL oscillator provides only an idealized description of dynamics near the Hopf bifurcation [3].
@@ 4 Discussion
Given the convincing agreement between experiments and simulations outlined above, it is worthwhile to further analyze the characteristic features of the amplitude patterns in more detail. The simulations confirm two remarkable properties of the patterns. First, although the systems are heterogeneous and the natural frequencies $\omega(\vec{x})$ are sampled from a Gaussian distribution, the measured frequencies $\Omega(\vec{x})$ in the bimodal pattern are spatially uniform. This indicates full frequency entrainment despite the imposed heterogeneity.
Second, under adiabatic continuation in the control parameter, the system does not exhibit parameter intervals with fixed size ratios $\rho$, as is typical for the globally coupled Stuart–Landau system (see [20]). Instead, the interface between the low- and high-amplitude domains shifts smoothly with the control parameter $C_{1}$, so that the size of the high-amplitude region, and hence $\rho$, changes continuously with $C_{1}$. In other words, rather than selecting one of several discrete (“quantized”) cluster ratios, the modified heterogeneous CGLE exhibits a near-continuum of bimodal patterns parameterized by the interface position. The same behavior is observed using the same random initial condition at each parameter set. However, when diffusion is removed, the system jumps randomly between coexisting cluster solutions. In the following, we will discuss the origin of these two peculiar properties of the patterns.
@@ 4.1 The Ghost of the cluster Singularity in Globally Coupled Heterogeneous Stuart-Landau oscillators
We first address the question of the mechanism that leads to complete frequency entrainment of the heterogeneous medium with Gaussian-distributed frequencies $\omega(\vec{x})$ even at moderate coupling strengths. This differs from phase-only models, such as the Kuramoto model, in which complete synchronization is only observed when the coupling strength $K$ is at least equal to the frequency deviation between the fastest (or slowest) oscillator and the mean frequency [4]. This suggests a crucial role of the amplitude degree of freedom. To understand the entrainment mechanism in amplitude oscillators, this section disregards the spatial extension, i.e., the diffusion term in Eq. (5), and studies an ensemble of only globally coupled oscillators, while retaining both amplitude and phase dynamics.
@@
$\partial_{t}W_{k}$ $=W_{k}(1-i\omega_{k})-(1+iC_{2})\lvert W_{k}\rvert^{2}W_{k}+K(1+iC_{1})\bigl(\langle W\rangle-W_{k}\bigr),$ (7)
@@
In the phase-locked regime, each oscillator can be written in polar form
@@
Returning to the frequency-locked amplitude mode, it would be desirable to study its emergence with linear stability analysis. However, this is not feasible in this case: the locked state is only implicitly defined through the self-consistency equations for $\Omega$ and $\mathfrak{R}$, which could not be obtained analytically. As a result, we focus on a qualitative and numerical analysis of how the system transitions from a unimodal to a bimodal amplitude distribution. Therefore, we analyze the fixed-point equations (11). By eliminating the phase variable $\phi_{k}$ via the identity $\sin^{2}\phi_{k}+\cos^{2}\phi_{k}=1$, we obtain an implicit equation for the amplitude $r$, writing $r_{k}\to r$ and $\phi_{k}\to\phi$ for clarity since we consider a single oscillator at fixed $\omega$:
@@
Due to the implicit dependence of $F(\vec{p})$ on $\mathfrak{R}$ and $\Omega$, we cannot continue the solution branches of the full system. However, we can numerically elucidate some properties of the transition from unimodal to bimodal states. Therefore, we fix $C_{1}=-2$, $C_{2}=2$, draw the intrinsic frequencies as $\omega_{k}\sim\mathcal{N}(0,0.1)$, and set $N=10^{4}$. We then numerically solve Eqs. (12) for coupling strengths $K\in\{1.5,1.3,1.2,1.0\}$. For each $K$, we determine the macroscopic parameters $\mathfrak{R}$ and $\Omega$ from simulations, and evaluate $F(\vec{p})$ for a range of $\omega$-values.
@@
Figure 7: Numerical simulation of system 7 for $N=10000$. Top row: amplitude distributions for different values of $K$. Bottom row: $F(\vec{p})$ for five different values of $\omega$. The same initial condition has been chosen for all simulations. The eigenfrequencies are pulled out of $\mathcal{N}(0,0.1)$ with the seed (MersenneTwister(2024). Other parameters are $C_{2}=2$ and $C_{1}=-2$.
@@
Next, let us compare the dynamics of the heterogeneous SL oscillators described above with the known dynamics of identical SL oscillators. In the parameter range considered here, identical and globally coupled SL oscillators exhibit amplitude clusters, consisting of two internally synchronized groups that show a constant phase difference; each group has its own amplitude [20, 21]. It therefore stands to reason that the bimodal amplitude distributions are the analog of 2-cluster solutions in heterogeneous oscillators.
This assignment calls for further comparison: In the case of identical oscillators, the full $\mathbf{S}_{N}$ permutation symmetry of the dynamical system leads to a huge multistability [19] between different cluster states, particularly in the region where the homogeneous solution is unstable. Clearly, the heterogeneous system lacks a permutation symmetry. However, the bistable states of the individual oscillators discussed above allow for many different amplitude distributions to satisfy Eq. (12). We can thus view the coexistence of other distributions as the inheritance of an identical oscillator system and speculate that the multiplicity is greater the narrower the frequency distribution.
Furthermore, in the homogeneous case, all the different 2-cluster states are born in saddle-node bifurcations which meet in a codimension-2 bifurcation, the cluster singularity [20, 21]. The unfolding of the cluster singularity reveals an ordering of the 2-clusters, where the most unbalanced clusters are born at the borders of the 2-cluster region, and successively more balanced ones are formed as one moves toward the center of the region. This results in a characteristic fan-like structure of the most probable cluster solution, as shown in Fig. 2 of Ref. [23]. There are indications that relics of the cluster singularity also exist in our heterogeneous mean-field system. This concerns the saddle-node bifurcations that generate the distributions split into two parts. Moreover, we observed that the relative ratio of the amplitude modes $\rho$ tends to increase with increasing $C_{1}$, reminiscent of the fan-like structure.
@@ 4.2 Pattern Selection Mechanism
We first assume the oscillators to be identical, and set $\omega(\vec{x})=\omega_{0}=0$. Furthermore, we consider the limit of small (non-zero) diffusion, $\epsilon\rightarrow 0$, in which the spatial profile of the amplitudes and phases closely resembles a step function. This is justified by the numerical results, where diffusion layers between clusters remain narrow even for moderate values of $\epsilon$ (cf. Fig.4) and allows us to adopt an analytical approach inspired by Refs. [43, 21]. To this end, we project the dynamics of the identical oscillators onto the center manifold of the Benjamin-Feir instability. This removes the neutral direction associated with phase invariance and reduces the dynamics of each oscillator to a scalar amplitude equation. The resulting one-component PDE reads:
@@
where $\lambda$ is a function of $K\text{ and }C_{2}$ and $A,B\text{ and},C$ are functions of $K,C_{1}\text{ and }C_{2}$ of our full Eq. (5[21]. The diffusion coefficient has not been subjected to the coordinate transformation, and we have added it to capture spatial structures. The variable $u(x,t)$ is thus a scalar function of space $x\in[0,L]$ and time, where we assume $x$ to be 1-dimensional. This scalar function can be treated as the amplitude of the Stuart–Landau oscillator. The global coupling is defined as:
@@
Moreover, the center manifold reduction imposes the constraint $\langle u\rangle=0$ [21], which yields:
@@
The above expression allows for two two-cluster fixed-point solutions [21], where each fixed-point solution expresses the amplitude of one of the clusters of the two-cluster solution. We note the two fixed points $u_{1}^{a,b}$. Notice that the sign of $u_{1}^{a}$ is always opposite to that of $u_{2}^{a}$. The expression of both fixed points is given by:
@@
where $u_{1}^{a}$ corresponds to the stable branch [21].
@@
To assess the predictive accuracy of the Lyapunov functional $\mathcal{F}(\rho,u)$ in capturing the selected cluster ratio $\rho$, we numerically solve equation (13) in one spatial dimension with Neumann boundary conditions for 800 oscillators and $L=10$. Simulations are initialized using a sigmoidal initial condition satisfying $\langle u\rangle=0$, and parameters are chosen in the vicinity of the cluster singularity (CS). Following Ref. [21], we fix
@@
The path $\mathcal{C}$ begins and ends at saddle-node bifurcations of solitary states, whose location in parameter space can easily be derived from Ref. [21]:
@@
In addition to cluster selection, diffusion also appears to define a critical domain size beyond which clustered areas do not exist. In both experiments and simulations, the ratio of high- and low-amplitude domains could only be tuned within an intermediate region. Close to $\rho=0$ and $\rho=1$, the system spontaneously returned to a unimodal distribution. Such a threshold is absent in discrete systems without diffusion. The synchronous solution coexists with cluster states at ratios close to 0 or 1, while diffusion naturally biases the system toward homogeneity. Two mechanisms may account for this transition: (i) the finite width of the diffusion layer, set by the ratio of the characteristic time scales of diffusion and reaction, destroys clusters smaller than this scale; and (ii) in two-dimensional media, the larger diffusion flux across curved fronts requires a critical nucleus size, below which domains collapse and the system reverts to a homogeneous profile [44, 45].
@@
At least phenomenologically, there appears to be a certain similarity between the amplitude cluster patterns described here and stationary domain patterns in bistable media with negative global coupling or feedback. In those systems, individual elements are bistable, and within a certain parameter range, a global constraint (e.g., a fixed mean field or imposed current) prevents the complete expansion of one state at the expense of the other, thereby stabilizing domain coexistence. By adjusting a parameter that impacts the global coupling, the size ratio of the two domains can usually be tuned, similar to what we describe above for the relative size of the cluster domains. Examples of such bistable systems include lithium-ion batteries, where at intermediate charging states, charged and discharged nanoparticles coexist, rather than a state composed of particles with an identical partial charging state [46]. Other examples include hot-wire barretters, which can serve as voltage stabilizers [47, 44] or the electrochemical oxidation of CO on Pt under galvanostatic conditions [48, 49, 50]. A notable difference is that all individual, uncoupled oscillators have a unique amplitude, with a value that differs from the amplitude values of both domains. This implies that the existence of different states is coupling-induced, whereas in bistable systems, the states are determined by the dynamics of the individual elements, with the global coupling being responsible for their stable coexistence. In the oscillatory system, cluster or domain formation requires a finite value of $C_{1}$ in Eq. (5). The variable $C_{1}$ translates to a cross-coupling of two variables in a physical system. In the one-component normal form 13, the cross-coupling manifests itself in the nonlinear coupling terms. In contrast, the global coupling in the one-component bistable system can be linear, the necessary nonlinearity for coexisting domains stemming from the dynamics of the individual elements. It thus seems that in the case of amplitude clusters, the complexity in the evolution equations is shifted from the individual element to the coupling. In what ways these two methods for generating domain coexistence can be mapped to each other remains an interesting open question.
@@ 5 Conclusion & Outlook
Neglecting diffusion and studying heterogeneous mean-field coupled SL oscillators revealed that the frequency heterogeneities are absorbed in the amplitude degree of freedom  [^1] The same qualitative effect occurs if heterogeneity is introduced via parameters other than $\omega_{k}$, such as the shear parameter $C_{2,k}$, which allows complete synchronization. The transition of the amplitude distributions from unimodal to bimodal could be related to the occurrence of saddle-node bifurcations in oscillators with a certain interval of heterogeneous frequencies, leading to a high degree of multistability, which can be easily verified numerically. A comparison with the dynamics of homogeneous mean-field coupled oscillators revealed that the organizing structure imposed by the cluster singularity on the two–cluster states in homogeneous Stuart–Landau ensembles persists in the presence of heterogeneity and diffusion. In other words, the ghost of the cluster singularity, which is a genuine property of systems with $\mathbf{S}_{N}$-symmetry, is still felt in the presence of small heterogeneities and diffusion.
@@
In short, our results show that synchronization in amplitude oscillators is facilitated by absorbing the heterogeneous eigendynamics in amplitude variations, and diffusion in mean-field coupled oscillatory media drives the highly multistable system to a unique state. Finally, we compared the mechanisms generating amplitude clusters in oscillatory media with the formation of two groups or domains in bistable systems with negative global coupling, which seem to share many properties.
This work raises several open questions. On the experimental side, further investigations are needed to clarify why the external resistance in the electrochemical system effectively acts as a cross-coupling term. Another question concerns stochastic perturbations, which are present in any experiment and may influence the stability and selection of patterns, but were not considered in this study. Moreover, it would be important to analyze the transition into incoherence for the heterogeneous system of globally coupled SL oscillators (7). Numerical evidence suggests that there is a discontinuous transition off frequency synchronization as $K$ is decreased past a critical parameter $K_{c}$. In this context, it is interesting to investigate how far it is related to more recently reported synchronization transitions.

### v2

@@ 3.2 Comparison of experimental and numerical results
Despite this agreement, systematic differences remain. Experimentally, high-amplitude regions pin to edges and form anisotropic shapes, while in simulations, they nucleate as compact patches. Such discrepancies arise from inhomogeneities, gradients, and boundary effects absent in the model, which assumes randomly distributed parameters (frequencies) and Neumann boundary conditions. Quantitative amplitude, phase, and frequency deviations are also natural, since the SL oscillator provides only an idealized description of dynamics near the Hopf bifurcation [Haken.1984].
@@ 4 Discussion
Given the convincing agreement between experiments and simulations outlined above, it is worthwhile to further analyze the characteristic features of the numerically obtained amplitude patterns in more detail.
The simulations confirm two remarkable properties of the patterns. First, although the systems are heterogeneous and the natural frequencies $\omega(\vec{x})$ are sampled from a Gaussian distribution, the measured frequencies $\Omega(\vec{x})$ in the bimodal pattern are spatially uniform. This indicates full frequency entrainment despite the imposed heterogeneity. Second, under adiabatic continuation in the control parameter, the system does not exhibit parameter intervals with fixed size ratios $\rho$, as is typical for the globally coupled Stuart–Landau system (see [Kemeth.2019]). Instead, the interface between the low- and high-amplitude domains shifts smoothly with the control parameter $C_{1}$, so that the size of the high-amplitude region, and hence $\rho$, changes continuously with $C_{1}$. In other words, rather than selecting one of several discrete (“quantized”) cluster ratios, the modified heterogeneous CGLE exhibits a near-continuum of bimodal patterns parameterized by the interface position. The same behavior is observed using the same random initial condition at each parameter set. However, when diffusion is removed, the system jumps randomly between coexisting cluster solutions. In the following, we will discuss the origin of these two peculiar properties of the patterns.
@@ 4.1 The Ghost of the Cluster Singularity in Globally Coupled Heterogeneous Stuart-Landau oscillators
We first address the question of the mechanism that leads to complete frequency entrainment of the heterogeneous medium with Gaussian-distributed frequencies $\omega(\vec{x})$ even at moderate coupling strengths. This differs from phase-only models, such as the Kuramoto model, in which complete synchronization is only observed when the coupling strength $K$ is at least equal to the frequency deviation between the fastest (or slowest) oscillator and the mean frequency [Strogatz.2000]. This suggests a crucial role of the amplitude degree of freedom. To understand the entrainment mechanism in amplitude oscillators, this section disregards the spatial extension, i.e., the diffusion term in Eq. (5), and studies an ensemble of only globally coupled Stuart-Landau oscillators. This complies with experimental conditions where the global coupling strength is much stronger than the diffusive coupling.
@@
$\partial_{t}W_{k}$ $=W_{k}(1-i\omega_{k})-(1+iC_{2})\lvert W_{k}\rvert^{2}W_{k}$ (8)
$+K(1+iC_{1})\bigl(\langle W\rangle-W_{k}\bigr),$
@@
In both experiments and simulations, we observed phase-locked solutions, which can be written in polar form
@@
Returning to the frequency-locked amplitude mode, introduced in equations (10), it would be desirable to study its emergence with linear stability analysis. However, this is not feasible in this case: the locked state is only implicitly defined through the self-consistency equations for $\Omega$ and $\mathfrak{R}$, which could not be obtained analytically. As a result, we focus on a qualitative and numerical analysis of how the system transitions from a unimodal to a bimodal amplitude distribution. Therefore, we analyze the fixed-point equations (12). By eliminating the phase variable $\phi_{k}$ via the identity $\sin^{2}\phi_{k}+\cos^{2}\phi_{k}=1$, we obtain an implicit equation for the amplitude $r$, writing $r_{k}\to r$ and $\phi_{k}\to\phi$ for clarity since we consider a single oscillator at fixed $\omega$:
@@
Due to the implicit dependence of $F(\vec{p})$ on $\mathfrak{R}$ and $\Omega$, we cannot continue the solution branches of the full system. However, we can numerically elucidate some properties of the transition from unimodal to bimodal states. Therefore, we fix $C_{1}=-2$, $C_{2}=2$, draw the intrinsic frequencies as $\omega_{k}\sim\mathcal{N}(0,0.1)$, and set $N=10^{4}$. We then numerically solve Eqs. (8) for coupling strengths $K\in\{1.5,1.3,1.2,1.0\}$. For each $K$, we determine the macroscopic parameters $\mathfrak{R}$ and $\Omega$ of the endstate of the simulation, and evaluate $F(\vec{p})$ for a range of $\omega$-values.
@@
Figure 7: Numerical simulation of system (8) for $N=10000$. Top row: amplitude distributions for different values of $K$. Bottom row: $F(\vec{p})$ for five different values of $\omega$ based on the parameters obtained from the simulations above. The same initial condition has been chosen for all simulations. The eigenfrequencies are pulled out of $\mathcal{N}(0,0.1)$ with the seed (MersenneTwister(2024). Other parameters are $C_{2}=2$ and $C_{1}=-2$.
@@
Next, let us compare the dynamics of the heterogeneous SL oscillators described above with the known dynamics of identical SL oscillators. In the parameter range considered here, identical and globally coupled SL oscillators exhibit amplitude clusters, consisting of two internally synchronized groups that show a constant phase difference; each group has its own amplitude [Kemeth.2019, Kemeth.2020]. It therefore stands to reason that the bimodal amplitude distributions are the analog of 2-cluster solutions in heterogeneous oscillators.
This assignment calls for further comparison: In the case of identical oscillators, the full $\mathbf{S}_{N}$ permutation symmetry of the dynamical system leads to a huge multistability [Golubitsky.2002] between different cluster states, particularly in the region where the homogeneous solution is unstable. Clearly, the heterogeneous system lacks a permutation symmetry. However, the bistable states of the individual oscillators discussed above allow for many different amplitude distributions to satisfy Eq. (13). We can thus view the coexistence of other distributions as the inheritance of an identical oscillator system and speculate that the multiplicity is greater the narrower the frequency distribution.
Furthermore, in the homogeneous case, all the different 2-cluster states are born in saddle-node bifurcations which meet in a codimension-2 bifurcation, the cluster singularity [Kemeth.2019, Kemeth.2020]. The unfolding of the cluster singularity reveals an ordering of the 2-clusters, where the most unbalanced clusters are born at the borders of the 2-cluster region, and successively more balanced ones are formed as one moves toward the center of the region. This results in a characteristic fan-like structure of the most probable cluster solution, as shown in Fig. 2 of Ref. [Thome.2025]. There are indications that relics of the cluster singularity also exist in our heterogeneous mean-field system. This concerns the saddle-node bifurcations that generate the distributions split into two parts. Moreover, we observed that the relative ratio of the amplitude modes $\rho$ tends to increase with increasing $C_{1}$, reminiscent of the fan-like structure.
@@ 4.2 Pattern Selection Mechanism
We first assume the oscillators to be identical, and set $\omega(\vec{x})=\omega_{0}=0$. Furthermore, we consider the limit of small (non-zero) diffusion, $\epsilon\rightarrow 0$, in which the spatial profile of the amplitudes and phases closely resembles a step function. This is justified by the numerical results, where diffusion layers between clusters remain narrow even for moderate values of $\epsilon$ (cf. Fig.4) and allows us to adopt an analytical approach inspired by Refs. [Dias.2003, Kemeth.2020]. To this end, we project the dynamics of the identical oscillators onto the center manifold of the Benjamin-Feir instability. This removes the neutral direction associated with phase invariance and reduces the dynamics of each oscillator to a scalar amplitude equation. The resulting one-component PDE reads:
@@
where $\lambda$ is a function of $K\text{ and }C_{2}$ and $A,B\text{ and},C$ are functions of $K,C_{1}\text{ and }C_{2}$ of our full Eq. (5[Kemeth.2020]. The diffusion coefficient has not been subjected to the coordinate transformation, and we have added it to capture spatial structures. The variable $u(x,t)$ is thus a scalar function of space $x\in[0,L]$ and time, where we assume $x$ to be 1-dimensional. This scalar function can be treated as the amplitude of the Stuart–Landau oscillator. The global coupling is defined as:
@@
Moreover, the center manifold reduction imposes the constraint $\langle u\rangle=0$ [Kemeth.2020], which yields:
@@
The above expression allows for two two-cluster fixed-point solutions [Kemeth.2020], where each fixed-point solution expresses the amplitude of one of the clusters of the two-cluster solution. We note the two fixed points $u_{1}^{a,b}$. Notice that the sign of $u_{1}^{a}$ is always opposite to that of $u_{2}^{a}$. The expression of both fixed points is given by:
@@
where $u_{1}^{a}$ corresponds to the stable branch [Kemeth.2020].
@@
To assess the predictive accuracy of the Lyapunov functional $\mathcal{F}(\rho,u)$ in capturing the selected cluster ratio $\rho$, we numerically solve equation (14) in one spatial dimension with Neumann boundary conditions for 800 oscillators and $L=10$. Simulations are initialized using a sigmoidal initial condition satisfying $\langle u\rangle=0$, and parameters are chosen in the vicinity of the cluster singularity (CS). Following Ref. [Kemeth.2020], we fix
@@
The path $\mathcal{C}$ begins and ends at saddle-node bifurcations of solitary states, whose location in parameter space can easily be derived from Ref. [Kemeth.2020]:
@@
In addition to cluster selection, diffusion also appears to define a critical domain size beyond which clustered areas do not exist: in both experiments and simulations, the ratio of high- and low-amplitude domains could only be tuned within an intermediate region, see Figures 3 and 6. Close to $\rho=0$ and $\rho=1$, the system spontaneously returned to a unimodal distribution. Such a threshold is absent in discrete systems without diffusion. The synchronous solution coexists with cluster states at ratios close to 0 or 1, while diffusion naturally biases the system toward homogeneity. Two mechanisms may account for this transition: (i) the finite width of the diffusion layer, set by the ratio of the characteristic time scales of diffusion and reaction, destroys clusters smaller than this scale; and (ii) in two-dimensional media, the larger diffusion flux across curved fronts requires a critical nucleus size, below which domains collapse and the system reverts to a homogeneous profile [Haken.1994, Haken.1996].
@@
At least phenomenologically, there appears to be a certain similarity between the amplitude cluster patterns described here and stationary domain patterns in bistable media with negative global coupling or feedback. In those systems, individual elements are bistable, and within a certain parameter range, a global constraint (e.g., a fixed mean field or imposed current) prevents the complete expansion of one state at the expense of the other, thereby stabilizing domain coexistence. By adjusting a parameter that impacts the global coupling, the size ratio of the two domains can usually be tuned, similar to what we describe above for the relative size of the cluster domains. Examples of such bistable systems include lithium-ion batteries, where at intermediate charging states, charged and discharged nanoparticles coexist, rather than a state composed of particles with an identical partial charging state [Dreyer.2011]. Other examples include hot-wire barretters, which can serve as voltage stabilizers [Barelko.1981, Haken.1994] or the electrochemical oxidation of CO on Pt under galvanostatic conditions [morschl.2008, bozdech.2018, Salman.2020]. A notable difference is that all individual, uncoupled oscillators have a unique amplitude, with a value that differs from the amplitude values of both domains. This implies that the existence of different states is coupling-induced, whereas in bistable systems, the states are determined by the dynamics of the individual elements, with the global coupling being responsible for their stable coexistence. In the oscillatory system, cluster or domain formation requires a finite value of $C_{1}$ in Eq. (5). The variable $C_{1}$ translates to a cross-coupling of two variables in a physical system. In the one-component normal form (14), the cross-coupling manifests itself in the nonlinear coupling terms. In contrast, the global coupling in the one-component bistable system can be linear, the necessary nonlinearity for coexisting domains stemming from the dynamics of the individual elements. It thus seems that in the case of amplitude clusters, the complexity in the evolution equations is shifted from the individual element to the coupling. In what ways these two methods for generating domain coexistence can be mapped to each other remains an interesting open question.
@@ 5 Conclusion & Outlook
Neglecting diffusion and studying heterogeneous mean-field coupled SL oscillators revealed that the frequency heterogeneities are absorbed in the amplitude degree of freedom, which allows complete synchronization  [^1] The same qualitative effect occurs if heterogeneity is introduced via parameters other than $\omega_{k}$, such as the shear parameter $C_{2,k}$.. The transition of the amplitude distributions from unimodal to bimodal could be related to the occurrence of saddle-node bifurcations in oscillators with a certain interval of heterogeneous frequencies, leading to a high degree of multistability, which can be easily verified numerically. A comparison with the dynamics of homogeneous mean-field coupled oscillators revealed that the organizing structure imposed by the cluster singularity on the two–cluster states in homogeneous Stuart–Landau ensembles persists in the presence of heterogeneity and diffusion. In other words, the ghost of the cluster singularity, which is a genuine property of systems with $\mathbf{S}_{N}$-symmetry, is still felt in the presence of small heterogeneities and diffusion.
@@
In short, our results show that synchronization in amplitude oscillators is facilitated by absorbing the heterogeneous eigendynamics in amplitude variations, and diffusion in mean-field coupled oscillatory media drives the highly multistable system to a unique state, in accordance with recent works [Millan.2025]. Finally, we compared the mechanisms generating amplitude clusters in oscillatory media with the formation of two groups or domains in bistable systems with negative global coupling, which seem to share many properties.
This work raises several open questions. On the experimental side, further investigations are needed to clarify why the external resistance in the electrochemical system effectively acts as a cross-coupling term. Another question concerns stochastic perturbations, which are present in any experiment and may influence the stability and selection of patterns, but were not considered in this study. Moreover, it would be important to analyze the transition into incoherence for the heterogeneous system of globally coupled SL oscillators (8). Numerical evidence suggests that there is a discontinuous transition off frequency synchronization as $K$ is decreased past a critical parameter $K_{c}$. In this context, it is interesting to investigate how far it is related to more recently reported synchronization transitions, such as explosive [GomezGardenes.2011] or extreme synchronization [Lee.2025].